a population has a mean of 200 and a standard deviation of 60. suppose a sample of size is 100 is selected and sample mean is used to estimate the mean. Use z table.
a. what is the probability that the sample mean will be within +/-7 of the population mean (to 4 decimals)
b. what is the probability that the sample mean will be within +/-16 of the population mean (to 4 decimals)
round z value in intermediate calculations to 2 decimal places
Answer #1

solution:-

given that mean = 200 , standard deviation = 60 , n = 100

a. the probability that the sample mean will be within +/- 7 of the population mean

z(193) = (193-200)/(60/sqrt(100)) = -1.17

z(207) = (207-200)/(60/sqrt(100)) = 1.17

=>p(-1.17 < z < 1.17) = p(z < 1.17) - p(z < -1.17)

= 0.8790 - 0.1210

= 0.7580


b. the probability that the sample mean will be within +/-16 of the population mean

z(184) = (184-200)/(60/sqrt(100)) = -2.67

z(216) = (216-200)/(60/sqrt(100)) = 2.67

P(-2.67 < z < 2.67) = p(z < 2.67) - p(z < -2.67)

= 0.9962 - 0.0038

= 0.9924